How do I solve for x in the first equation and for w in the next?
– 3/4x – 2 = - 8
and
5/3(9 + w) = - 10

Answer by Alan3354(69443)

How do I solve for x in the first equation and for w in the next?
– 3/4x – 2 = - 8
and
5/3(9 + w) = - 10
------------------
You get the variable isolated on one side of the equal sign.
– 3/4x – 2 = - 8
Start by adding 2 to both sides.
Make an effort.